OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. y – 4 = 2(x + 3) y + 4 = 2(x – 3) y + 3 = 2(x – 4) y – 3 = 2(x + 4)

OpenStudy (anonymous):

y-4=2(x+3)

OpenStudy (anonymous):

y-y1=m(x-x1) is the formula, then you just sub in your values

OpenStudy (anonymous):

can you do this other one please

OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

y -4 = 2 ( x +3)